# An Efficient Instance Segmentation Framework Based on Oriented Bounding Boxes

### v3

@@ Abstract
Instance segmentation for completely occluded objects and dense objects in robot vision measurement are two challenging tasks. To uniformly deal with them, this paper proposes a unified coarse-to-fine instance segmentation framework, CFNet, which uses box prompt-based segmentation foundation models (BSMs), e.g., Segment Anything Model. Specifically, CFNet first detects oriented bounding boxes (OBBs) to distinguish instances and provide coarse localization information. Then, it predicts OBB prompt-related masks for fine segmentation. CFNet performs instance segmentation with OBBs that only contain partial object boundaries on occluders to predict occluded object instances, which overcomes the difficulty of existing amodal instance segmentation methods in directly predicting occluded objects. In addition, since OBBs only serve as prompts, CFNet alleviates the over-dependence on bounding box detection performance of current instance segmentation methods using OBBs for dense objects. Moreover, to enable BSMs to handle OBB prompts, we propose a novel OBB prompt encoder. To make CFNet more lightweight, we perform knowledge distillation on it and introduce a Gaussian label smoothing method for teacher model outputs. Experiments demonstrate that CFNet outperforms current instance segmentation methods on both industrial and public datasets. The code is available at https://github.com/zhen6618/OBBInstanceSegmentation.
@@ I INTRODUCTION
Instance segmentation provides foundational information for numerous tasks based on robot vision measurement, such as robot grasping [1, 2] and autonomous driving [3]. This work focuses on two challenging difficulties in robot vision measurement: completely occluded object instance segmentation and dense object instance segmentation. We aim to develop a unified framework to handle these difficulties.
Figure 1: Examples of completely occluded objects (a) and dense objects (b). Original objects are inside the corresponding black dotted boxes. (a): reference holes (blue) are occluded by bolts or nuts (i.e., occluders, shown in red). Oriented bounding boxes (green) contain occluder boundaries (orange) that are located at the contact surface between occluders and occluded reference holes. (b): dense vehicles (red) are surrounded by horizontal bounding boxes (yellow) or oriented bounding boxes (green).
Instance segmentation for completely occluded objects (called occludees) is difficult, since the information available for inferring occluded objects is very limited. For example, in Fig. 1(a), instance segmentation of reference holes that provide optimal assembly positions is usually required in robot assembly tasks. In several scenarios, reference holes (i.e., occludees) are screwed with bolts and nuts, completely occluded by these industrial parts (called occluders). To perform instance segmentation on occludees, current amodal instance segmentation methods [4, 5, 6, 7] directly predict occludees. Although they can automatically extract features related to occludees for inference, the performance is poor.
@@
Dense object instance segmentation is a long-standing challenge [8] in robot vision measurement. For instance, in Fig. 1(b), objects are densely packed in multiple orientations in unmanned aerial vehicle (UAV) measurement. Since horizontal bounding boxes (HBBs) introduce many interference areas, oriented bounding boxes (OBBs) are often used for instance identification [9, 10, 11, 12], e.g., ISOP [12] predicts masks in the detected OBBs. However, bounding box-based instance segmentation methods are overly dependent on bounding box detection performance, especially for OBBs that are sensitive to both position and orientation. This dependence makes instance segmentation methods using OBBs more difficult and unstable.
@@
In this paper, we propose CFNet, a unified coarse-to-fine framework for both completely occluded and dense object instance segmentation, which uses box prompt-based segmentation foundation models (BSMs), e.g., Segment Anything Model (SAM) [13]. CFNet first detects OBBs to distinguish instances, identify classes, and provide coarse localization information. Then, it predicts OBB-related segmentation masks. Our motivation is inspired by the powerful segmentation capabilities of BSMs and their excellent box prompt mechanisms. CFNet uses an OBB detection module (e.g., Oriented R-CNN [14]) for coarse detection, and leverages BSMs to predict OBB-related masks for fine segmentation.
@@
CFNet improves BSMs in two aspects. First, to handle OBB prompts, CFNet introduces a novel OBB prompt encoder. Second, since BSMs usually have high computational complexity, we perform knowledge distillation on the box prompt encoders and mask decoders. Specifically, a Gaussian smoothing method for teacher model outputs is introduced. In this work, CFNet chooses SAM as the baseline model and makes improvements in the above two aspects.
Compared with amodal instance segmentation methods, CFNet avoids directly predicting occludee instances by performing instance segmentation with OBBs that only contain partial object boundaries on occluders to infer occludees. In addition, OBBs used in CFNet provide more accurate location information than HBBs used in amodal instance segmentation methods. Compared with current instance segmentation methods using OBBs for dense objects, CFNet only uses OBBs as prompts to guide object segmentation, so the segmentation results are less dependent on bounding box detection performance, which improves instance segmentation performance. Moreover, CFNet is based on BSMs that are pretrained on large-scale data for segmentation tasks, helping to enhance feature extraction and segmentation capabilities. The main contributions are summarized as follows.
  • We propose CFNet, a unified coarse-to-fine framework for both completely occluded and dense object instance segmentation in robot vision measurement. CFNet uses BSMs and outperforms current instance segmentation methods on both industrial and public datasets.
@@
  • We propose a knowledge distillation method for the OBB prompt encoders and mask decoders of BSMs, significantly reducing the computational complexity with a minor loss in accuracy. Specifically, a Gaussian smoothing method for teacher model outputs is proposed.
@@ II Related Work
II-A Amodal Instance Segmentation
Amodal instance segmentation directly predicts completely occluded object instances. The first amodal instance segmentation method [7] predicted the amodal masks and corresponding bounding boxes in an iterative bounding box expansion manner. Zhu et al. [15] directly trained general instance segmentation methods on amodal instance segmentation datasets. Based on Mask R-CNN [16], Qi et al. [5] predicted occluded parts by adding a multi-task branch with multi-level coding. Similarly, based on Mask R-CNN, ORCNN [6] first predicted amodal masks through an amodal mask prediction branch and then combined a visible mask prediction branch to predict occluded parts. Different from the above methods that only considered occludees, BCNet [4] proposed a bilayer graph convolutional network to simultaneously consider interactions between occluders and occludees. Since the information available for inferring occludees is very limited, these methods have poor performance.
II-B Instance Segmentation with Oriented Bounding Boxes
@@
Compared with instance segmentation with HBBs [16, 17], instance segmentation with OBBs that have less interference areas provides more accurate location information. Based on Mask R-CNN, a Region of Interest (RoI) learner was applied to HBB proposals to generate OBB proposals, followed by a head that generated segmentation masks [12]. Based on a similar idea, Feng et al. [11] proposed a part-aware instance segmentation network with OBB proposals for bin picking. Follmann and König [10] first generated the final OBBs using a two-stage object detection approach, and then predicted masks within the detected OBBs. Rotated Blend Mask R-CNN [9] proposed a top-down and bottom-up structure for oriented instance segmentation. However, these methods are overly dependent on OBB detection performance, making predictions more difficult and unstable.
II-C Box Prompt-Based Segmentation Foundation Models
@@
Benefiting from pretraining on large-scale data for segmentation tasks, BSMs exhibit powerful segmentation and generalization capabilities. For example, SAM [13] was trained on more than 1B segmentation masks from 11M images, showing remarkable segmentation and zero-shot generalization capabilities. A BSM for medical image segmentation is proposed in [18]. To enhance interactive performance, BSMs incorporate prompt mechanisms. A representative work is SAM, which took HBBs as prompts and predicted segmentation masks with respect to HBBs. MobileSAM [19] retained the HBB prompt mechanism and proposed a lightweight version of SAM for edge devices.
Figure 2: Architecture of the proposed CFNet. Compared with dense object instance segmentation, completely occluded object instance segmentation needs more post-processing steps to transform occluder instances into occludee instances.
II-D Box Prompt Encoder
@@
Existing methods usually use HBBs as box prompts for BSMs, such as SAM and MobileSAM. The current designs of HBB prompt encoders are mainly inspired by SAM. A HBB was represented by two points, i.e., the top-left corner point and the bottom-right corner point. Then, a point was encoded as the sum of a Gaussian positional encoding [20] of the location and a learned embedding representing the top-left corner or bottom-right corner. Hence, a box was encoded as a combination of the two encoded point embeddings. However, to the best of our knowledge, there is no relevant research on OBB prompt encoders yet.
II-E Label Smoothing for Box Prompt encoders and Segmentation networks in Knowledge Distillation
@@
In knowledge distillation for segmentation tasks, label smoothing helps reduce overfitting and enhances model performance. Inception-v3 [21] proposed uniform label smoothing (ULS), i.e., the weighted average of one-hot labels and uniform distributions. Spatially varying label smoothing (SVLS) [22] applied a discrete spatial Gaussian kernel to segmentation labels to smooth one-hot labels. These two label smoothing methods were commonly used in segmentation tasks. For example, Park et al. [23] applied ULS and proposed pixel-wise adaptive label smoothing for self-distillation. P-CD [24] utilized SVLS to generate segmentation boundary uncertainty and soft targets. Unlike applying SVLS to ground truth (GT) labels, we apply Gaussian smoothing to the segmentation masks of the teacher model, which yields better distillation performance. As for label smoothing methods used for box prompt encoders, to our knowledge, no relevant research has been published yet.
@@ III-A Overview
The structure of CFNet is shown in Fig. 2. It is mainly composed of four parts: an OBB detection module, an image encoder, an OBB prompt encoder, and a mask decoder. Specifically, images containing dense or occluded objects serve as input. The OBB detection module is used to detect OBBs that distinguish instances, identify classes, and provide coarse localization information. The image encoder is used to transform input images into high-dimensional feature representations and generate image embeddings. The OBB prompt encoder is responsible for encoding OBB prompts and generating prompt embeddings. Then, the image embeddings and prompt embeddings are combined in the mask decoder to predict OBB prompt-related segmentation masks.
For dense object instance segmentation, CFNet detects OBBs that encompass entire instances and predicts segmentation masks of dense objects. For completely occluded object instance segmentation, CFNet avoids directly predicting occludee instances by performing instance segmentation with OBBs that only contain partial object boundaries on occluders. Specifically, OBBs contain occluder boundaries that are located at the contact surface between occluders and occludees (see Fig. 1(a)). Then, post-processing steps transform occluder instances into partial boundary instances and use prior geometric relationships between boundary instances and occludee instances to obtain occludee instances (e.g., in robot assembly, the boundary shape of the occluded reference holes is an ellipse). However, partial bounding boxes are not suitable for current instance segmentation methods since their bounding boxes are usually designed to encompass entire instances.
Furthermore, knowledge distillation is applied to the OBB prompt encoder and mask decoder to reduce the computational complexity and make CFNet more lightweight. For OBB detection, we directly employ Oriented R-CNN [14], which is a two-stage OBB detection model with competitive detection accuracy and inference speed.
@@ III-B Oriented Bounding Box Prompt Encoder
Coordinate-based multilayer perceptrons (MLPs) take low-dimensional coordinates as input, such as $\bm{\varphi}_{1}$, $\bm{\varphi}_{2}$ and $\bm{\theta}$, and they have difficulty learning high-frequency information [20]. In CFNet, there are many MLP structures and transformer structures containing MLPs, so such a problem will degrade instance segmentation performance. To alleviate this problem, inspired by Gaussian positional encoding (GPE) for coordinates [20], CFNet introduces adaptive GPE (AGPE) to map the coordinate set ($\bm{\varphi}_{1},\bm{\varphi}_{2},\bm{\theta}$) into random Fourier features to better learn high-frequency information and enhance the performance of coordinate-based MLPs.
@@
In many HBB prompt encoders of BSMs, $(\bm{\rho}{\bm{\varphi}_{1}},\bm{\rho}{\bm{\varphi}_{2}},\bm{\rho}{\bm{\theta}})$ and ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$) are considered as hyperparameters, which leads to difficulty in adapting to changes in datasets and searching optimal parameter settings. In CFNet, we adaptively learn ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$). $(\bm{\rho}{\bm{\varphi}_{1}},\bm{\rho}{\bm{\varphi}_{2}},\bm{\rho}{\bm{\theta}})$ are fixedly set to 0. Concretely, ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$) are the weights of a learnable embedding layer.
Inspired by SAM, the mapped random Fourier features ($\gamma(\bm{\varphi}_{1})$ and $\gamma(\bm{\varphi}_{2})$) are summed with two learned embeddings ($\bm{\omega}_{1}$ and $\bm{\omega}_{2}$) that represent the top-left corner point and bottom-right corner point in CFNet, respectively.
@@
On the other hand, the parameterized orientation representation ($\sin\theta,\cos\theta$) of an OBB suffers from the boundary discontinuity problem [25]. Due to the periodicity of orientations, at boundary orientations (such as $0^{\circ}$ and $180^{\circ}$), small changes in orientations will result in large jumps in ($\sin\theta,\cos\theta$). To alleviate the boundary discontinuity problem, CFNet introduces learnable orientation correction embeddings $\bm{\omega}_{\theta}$. At the orientations where the boundary discontinuity problem occurs, $\bm{\omega}_{\theta}$ adaptively adjusts the effect of this problem. Hence, the encoded feature embeddings $\mathcal{E}(\bm{\theta})$ with respect to orientation information are

### v6

# OBSeg: Accurate and Fast Instance Segmentation Framework Using Segmentation Foundation Models with Oriented Bounding Box Prompts

@@ Abstract
Instance segmentation in remote sensing images is a long-standing challenge. Since horizontal bounding boxes introduce many interference objects, oriented bounding boxes (OBBs) are usually used for instance identification. However, based on “segmentation within bounding box” paradigm, current instance segmentation methods using OBBs are overly dependent on bounding box detection performance. To tackle this problem, this paper proposes OBSeg, an accurate and fast instance segmentation framework using OBBs. OBSeg is based on box prompt-based segmentation foundation models (BSMs), e.g., Segment Anything Model. Specifically, OBSeg first detects OBBs to distinguish instances and provide coarse localization information. Then, it predicts OBB prompt-related masks for fine segmentation. Since OBBs only serve as prompts, OBSeg alleviates the over-dependence on bounding box detection performance of current instance segmentation methods using OBBs. Thanks to OBB prompts, OBSeg outperforms other current BSM-based methods using HBBs. In addition, to enable BSMs to handle OBB prompts, we propose a novel OBB prompt encoder. To make OBSeg more lightweight and further improve the performance of lightweight distilled BSMs, a Gaussian smoothing-based knowledge distillation method is introduced. Experiments demonstrate that OBSeg outperforms current instance segmentation methods on multiple datasets in terms of instance segmentation accuracy and has competitive inference speed. The code is available at https://github.com/zhen6618/OBBInstanceSegmentation.
Index Terms:
Instance segmentation in remote sensing images, oriented bounding boxes, segmentation foundation models, knowledge distillation, oriented bounding box prompt encoder.
@@ I INTRODUCTION
Instance segmentation of remote sensing images provides pixel-level foundational information for numerous tasks, such as maritime security [1, 2], agricultural measurement [3], urban management, and disaster assessment. Pixel-level perception enables related instance segmentation methods to depict the exterior appearance and contour of the object. However, objects in remote sensing images usually have the characteristics of multi-scale, dense arrangement, and arbitrary orientations, which brings a long-standing challenge to the instance segmentation task.
@@
Since objects are densely packed in multiple orientations, which introduces many interference regions into horizontal bounding boxes (HBBs) [4, 5], oriented bounding boxes (OBBs) are usually used for instance identification [6, 7, 8, 9] (see Fig. 1). Existing instance segmentation methods are mainly based on the “segmentation within bounding box” paradigm, that is, object segmentation is mainly performed in the detected bounding box (ignoring small differences caused by RoIAlign). This paradigm makes the segmentation performance overly dependent on the bounding box detection performance. However, the core of the instance segmentation task is to distinguish different object instances and segment each instance with a corresponding mask. Bounding boxes can be used for instance identification and roughly locating instances, so bounding boxes can be regarded as tools to assist instance segmentation. If the segmentation is mainly restricted to the inside of the bounding boxes, the performance of instance segmentation will be degraded. Furthermore, since OBBs are sensitive to both position and orientation and are difficult to detect, this dependence makes instance segmentation methods using OBBs more challenging and unstable.
Figure 1: For instance segmentation in remote sensing images, (a): HBB introduces many interference objects. (b): The “segmentation within bounding box” paradigm limits the segmentation to be performed mainly within the detected OBB, making the segmentation performance overly dependent on the OBB detection performance. Once the OBB detection is inaccurate, the mask segmentation will also be affected. (c) The proposed OBSeg only uses OBB as a prompt to guide object segmentation, so the segmentation result is less dependent on OBB detection performance. Although the OBB detection is inaccurate, the mask can be segmented accurately.
@@
In this paper, to reduce dependence on bounding box detection performance, we propose OBSeg, an accurate and fast instance segmentation framework that uses box prompt-based segmentation foundation models (BSMs), e.g., Segment Anything Model (SAM) [10]. Our motivation is inspired by the powerful segmentation capabilities of BSMs and their excellent box prompt mechanisms. Although there are currently some instance segmentation methods [11] using BSMs (e.g., RSPrompter [11] improves on SAM), they use HBBs as prompts, introducing many interference regions. Different from BSMs with HBB prompts, OBSeg utilizes more accurate OBBs as prompts, which exhibits better segmentation performance. Specifically, OBSeg first detects OBBs to distinguish instances, identify classes, and provide coarse localization information. Then, it predicts OBB-related segmentation masks. OBSeg uses an OBB detection module [12, 13, 14] for coarse detection, and leverages BSMs to predict OBB-related masks for fine segmentation.
@@
OBSeg improves BSMs in two aspects. First, to handle OBB prompts, OBSeg introduces a novel OBB prompt encoder. Second, since BSMs usually have high computational complexity, we perform knowledge distillation [15] on BSMs to improve the practicality. Although there are many methods [16, 17, 18] that use knowledge distillation to introduce lightweight BSMs, the performance gap between the distilled model and the teacher model needs to be further reduced. In OBSeg, a Gaussian smoothing method for teacher model outputs is introduced to further improve distillation performance.
Compared with the “segmentation within bounding box” paradigm, OBSeg only uses OBBs as prompts to guide object segmentation, so the segmentation results are less dependent on bounding box detection performance, which effectively improves instance segmentation performance. Moreover, after distillation, OBSeg is more lightweight. The main contributions are summarized as follows.
  1. 1.
    We propose OBSeg, an accurate and fast instance segmentation framework using BSMs with OBB prompts. OBSeg outperforms current instance segmentation methods on multiple datasets in terms of instance segmentation accuracy and has competitive inference speed.
  2. 2.
@@
  3. 3.
    A Gaussian smoothing-based knowledge distillation method is introduced to improve the instance segmentation performance of lightweight distilled BSMs.
@@ II Related Work
II-A Instance Segmentation with Oriented Bounding Boxes
@@
Compared with instance segmentation methods using HBBs [19, 20], instance segmentation methods using OBBs that have fewer interference regions provide more accurate location information. Most of existing methods build upon the “segmentation within bounding box” paradigm and the improved Mask R-CNN [19] framework. For example, a Region of Interest (RoI) learner was applied to HBB proposals to generate OBB proposals in ISOP [9], followed by a head that generated segmentation masks for OBB proposals. Follmann and König [7] first generated the final OBBs using a two-stage object detection approach, and then predicted masks within the detected OBBs. Rotated Blend Mask R-CNN [6] proposed a top-down and bottom-up structure for oriented instance segmentation. Although the above methods effectively improved current instance segmentation performance in remote sensing images, they are overly dependent on OBB detection performance, making predictions more difficult and unstable. Our proposed OBSeg uses OBBs as prompts to guide object segmentation, which effectively reduces this dependence.
II-B Box Prompt-Based Segmentation Foundation Models
@@
Benefiting from pretraining on large-scale data for segmentation tasks, BSMs exhibit powerful segmentation and generalization capabilities. For example, SAM [10] was trained on more than 1B segmentation masks from 11M images, showing remarkable segmentation and zero-shot generalization capabilities. A BSM for medical image segmentation is proposed in [21]. To enhance interactive performance, BSMs introduce prompt mechanisms. A representative work is SAM, which took HBBs as prompts and predicted segmentation masks with respect to HBBs. To improve the practicality of BSMs, EfficientSAM [16] and MobileSAM [17] used knowledge distillation to introduce lightweight image encoders for edge devices. EdgeSAM [18] performs distillation on the image encoder, prompt encoder, and mask decoder. However, the performance of these distilled BSMs is limited. We introduce a Gaussian smoothing method for teacher model outputs to further improve distillation performance for BSMs.
Figure 2: Architecture of the proposed OBSeg. It is mainly composed of four parts: an OBB detection module, an image encoder, an OBB prompt encoder, and a mask decoder. OBSeg first detects OBBs to distinguish instances, identify classes, and provide coarse localization information. Then, the mask decoder utilizes the image embeddings generated by the image encoder and the OBB prompt embeddings generated by the OBB prompt encoder to generate segmentation masks. In addition, Gaussian smoothing-based knowledge distillation is performed on the OBB prompt encoder and the mask decoder to make OBSeg more lightweight.
II-C Box Prompt Encoder
@@
Existing methods usually use HBBs as box prompts for BSMs, such as SAM and EfficientSAM. The current designs of HBB prompt encoders are mainly inspired by SAM. A HBB was represented by two points, i.e., the top-left corner point and the bottom-right corner point. Then, a point was encoded as the sum of a Gaussian positional encoding [22] of the location and a learned embedding representing the top-left corner or bottom-right corner. Hence, a box was encoded as a combination of the two encoded point embeddings. In remote sensing images, using OBBs helps improve instance segmentation performance. However, to the best of our knowledge, there is no relevant research on OBB prompt encoders yet, which is necessary for BSMs with OBB prompts.
II-D Label Smoothing for Segmentation Networks in Knowledge Distillation
@@
In knowledge distillation for segmentation tasks, label smoothing helps reduce overfitting and enhances model performance. Inception-v3 [23] proposed uniform label smoothing (ULS), i.e., the weighted average of one-hot labels and uniform distributions. Spatially varying label smoothing (SVLS) [24] applied a discrete spatial Gaussian kernel to segmentation labels to smooth one-hot labels. These two label smoothing methods were commonly used in segmentation tasks. For example, Park et al. [25] applied ULS and proposed pixel-wise adaptive label smoothing for self-distillation. P-CD [26] utilized SVLS to generate segmentation boundary uncertainty and soft targets. Different from applying SVLS to ground truth (GT) labels, to further improve the generalization capability and instance segmentation performance of the student model, we apply Gaussian smoothing to the segmentation masks of the teacher model, which yields better distillation performance.
@@ III-A Overview
The structure of OBSeg is shown in Fig. 2. It is mainly composed of four parts: an OBB detection module, an image encoder, an OBB prompt encoder, and a mask decoder. Specifically, the OBB detection module is used to detect OBBs that distinguish instances, identify classes, and provide coarse localization information. The image encoder is used to transform input images into high-dimensional feature representations and generate image embeddings. The OBB prompt encoder is responsible for encoding OBB prompts and generating prompt embeddings. Then, the image embeddings and prompt embeddings are combined in the mask decoder to predict OBB prompt-related segmentation masks. Furthermore, Gaussian smoothing-based knowledge distillation is applied to the OBB prompt encoder and mask decoder to make OBSeg more lightweight and practical.
@@ III-B Oriented Bounding Box Prompt Encoder
Coordinate-based multilayer perceptrons (MLPs) take low-dimensional coordinates as input, such as $\bm{\varphi}_{1}$, $\bm{\varphi}_{2}$ and $\bm{\theta}$, and they have difficulty learning high-frequency information [22]. In OBSeg, there are many MLP structures and transformer structures containing MLPs, so such a problem will degrade instance segmentation performance. To alleviate this problem, inspired by Gaussian positional encoding (GPE) for coordinates [22], OBSeg introduces adaptive GPE (AGPE) to map the coordinate set ($\bm{\varphi}_{1},\bm{\varphi}_{2},\bm{\theta}$) into random Fourier features to better learn high-frequency information and enhance the performance of coordinate-based MLPs.
@@
In many HBB prompt encoders of BSMs, $(\bm{\rho}{\bm{\varphi}_{1}},\bm{\rho}{\bm{\varphi}_{2}},\bm{\rho}{\bm{\theta}})$ and ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$) are considered as hyperparameters, which leads to difficulty in adapting to changes in datasets and searching optimal parameter settings. In OBSeg, we adaptively learn ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$). $(\bm{\rho}{\bm{\varphi}_{1}},\bm{\rho}{\bm{\varphi}_{2}},\bm{\rho}{\bm{\theta}})$ are fixedly set to 0. Concretely, ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$) are the weights of a learnable embedding layer.
Inspired by SAM, the mapped random Fourier features ($\gamma(\bm{\varphi}_{1})$ and $\gamma(\bm{\varphi}_{2})$) are summed with two learned embeddings ($\bm{\omega}_{1}$ and $\bm{\omega}_{2}$) that represent the top-left corner point and bottom-right corner point in OBSeg, respectively.
@@
On the other hand, the parameterized orientation representation ($\sin\theta,\cos\theta$) of an OBB suffers from the boundary discontinuity problem [27]. Due to the periodicity of orientations, at boundary orientations (such as $0^{\circ}$ and $180^{\circ}$), small changes in orientations will result in large jumps in ($\sin\theta,\cos\theta$). To alleviate the boundary discontinuity problem, OBSeg introduces learnable orientation correction embeddings $\bm{\omega}_{\theta}$. At the orientations where the boundary discontinuity problem occurs, $\bm{\omega}_{\theta}$ adaptively adjusts the effect of this problem. Hence, the encoded feature embeddings $\mathcal{E}(\bm{\theta})$ with respect to orientation information are